# Quantitative U/Th deposition and cleanliness control strategies in the JUNO site air

### v1

@@ VI.2 LS purification and filling systems
The JUNO LS is subjected to a sequential four-stage purification process, which includes alumina adsorption, distillation, water extraction, and gas stripping. After this process, it is filled into the acrylic sphere by a filling system. The primary objectives of the alumina and gas stripping systems are to enhance the transparency of the alkylbenzene feedstock and to remove gaseous components from the LS, respectively. Consequently, their capabilities for the purification of U/Th are somewhat limited. The efficiency of the distillation system Landini:2024ohd and water extraction system Ye:2021jkp in removing U/Th impurities is expected to be 99% and 80% respectively. The total inner surface area of the entire system, as well as the surface area weighted by purification efficiency, are summarized in Table 6. With U/Th contamination in the 20 kt LS reaching levels of 10-18 g/g, corresponding to a U/Th mass of 20 ng. Therefore, for the 1105 m2 inner surface area of the LS-related system, the required U/Th residual should be smaller than 18 pg/m2. The U/Th settling rate in class 10,000 environment is estimated at about 400 pg/d/m2 based on the results of No.2 PFA bottle facing upwards in Table 3, allowing for an exposure time of 65 min in class 10,000 environment. During construction, all interfaces of pipelines and containers are covered with blind plates before docking, and the blind plates are removed during docking. The vast majority of interfaces strictly require an air exposure time during docking of less than 10 min, with internal nitrogen gas protection after docking is completed. Therefore, even if the environmental cleanliness level fluctuates several times higher than the class 10,000 level, the impact of air exposure during the installation of the LS-related system is controllable.
@@ VI.3 Structure outside the acrylic
The installation situation and progress for acrylic sphere, PMT and SS truss are summarized in Table 7, as well as the expected U/Th settling rate. Details on the construction can be found in Section IV.2. The exposure here only include the time after the final cleaning and before filling. The U/Th settling used in the calculation is the results of PFA flask from Table 4 during the exposure with the sphere facing upwards (No. 5 for acrylic and truss, No. 6 for PMT layer). The single-event rate within LS from the settled U/Th on the detector surface, and the release of radon in water are listed in the last two rows of the table. Overall, the U/Th contamination introduced during the installation of the detector due to environmental exposure contributes approximately 2 mHz to the single-event rate in the LS fiducial volume (R 17.2 m), accounting for about 0.03% of the total detector contribution (7 Hz JUNO:2021kxb ). The impact of radon released from the decay of 238U deposited on the detector surface into the water is at a level of 2 mBq/m3 by assuming equilibrium rate as 238U, which is approximately 20% of the water radon target (10 mBq/m3). Additionally, some of the deposited dust can be further purified through the water circulation process.
Acrylic | PMT | SS truss | |
@@ VII Summary
The installation of the JUNO detector spanned three years. To ensure the detector’s cleanliness, the on-site installation environment was managed under clean-room conditions, with an average cleanliness level of class 74,000. After the installation of the acrylic sphere and before the filling with LS, the first step of the cleaning involved a water mist dust reduction inside the sphere, which improved the air cleanliness to a level of class 100 to 1,000. The second step involved using a 3D high-pressure nozzle to rinse the inner surface of the acrylic sphere, further enhancing the cleanliness of the surface in contact with the LS. Surface deposition experiments for U/Th on various containers were conducted in the JUNO experimental hall. Based on the experimental results, the total U/Th deposition during the installation process of the entire detector was evaluated. Through detector simulations, it was determined that the contribution to the single-event rate in the LS fiducial volume is approximately 0.03% of the total rate from detector materials (7 Hz), and the radon contamination in water is at a level of 2 mBq/m3, which is 20% of the requirement (10 mBq/m3).
@@ Acknowledgements
This work is supported by the Youth Innovation Promotion Association of the Chinese Academy of Sciences, CAS Project for Young Scientists in Basic Research (YSBR-099).

### v2

@@ VI.2 LS purification and filling systems
The JUNO LS is subjected to a sequential four-stage purification process, which includes alumina adsorption, distillation, water extraction, and gas stripping. After this process, it is filled into the acrylic sphere by a filling system. The primary objectives of the alumina and gas stripping systems are to enhance the transparency of the alkylbenzene feedstock and to remove gaseous components from the LS, respectively. Consequently, their capabilities for the purification of U/Th are somewhat limited. The efficiency of the distillation system Landini and others (2024) and water extraction system Ye and others (2022) in removing U/Th impurities is expected to be 99% and 80% respectively. The total inner surface area of the entire system, as well as the surface area weighted by purification efficiency, are summarized in Table 5. With U/Th contamination in the 20 kt LS reaching levels of 10-18 g/g, corresponding to a U/Th mass of 20 ng. Therefore, for the 1105 m2 inner surface area of the LS-related system, the required U/Th residual should be smaller than 18 pg/m2.
Most pipes, containers, and components in the LS-related systems that are acid-resistant underwent degreasing, acid pickling, passivation, and UPW rinsing. The final validation of cleanliness included the following criteria:
The particle concentration in the rinse water must meet Level 50 standard (MIL-STD-1246C).
The change in resistivity before and after UPW rinsing should remain within 4–6 Mcm.
There is no significant change in the absorption spectrum of the rinse water before and after rinsing.
U/Th concentrations in the rinse water must be below 10-13 g/g.
Only when all these criteria are met is the component considered qualified. Similar to the challenges in acrylic surface cleaning, the exact amount of residual contamination on cleaned surfaces remains unknown. Therefore, we implemented multiple verification methods to ensure the effectiveness of the cleaning process.
After cleaning, all these components were sealed with nitrogen. During installation, however, these systems may be exposed to ambient air for limited periods. The acceptable duration of such exposure can be estimated based on the measurement results presented in the previous section. The cleanliness of the room housing the underground LS purification and filling systems has also been monitored long-term, maintaining a class 1,000-10,000 standard. These systems are mostly cylindrical, so we plan to use the measurement results from PFA bottles for background evaluation. Based on the results from PFA bottle No.2 (facing upwards) in Table 2, the U/Th settling rate in a Class 10,000 environment is estimated to be approximately 100 pg/d/m², scaled according to the cleanliness level. This estimate accounts for an exposure time of 4.3 hours under Class 10,000 conditions. During construction, all interfaces of pipelines and containers are covered with blind plates before docking, and the blind plates are removed during docking. The vast majority of interfaces strictly require an air exposure time during docking of less than 10 min, with internal nitrogen gas protection after docking is completed. Therefore, even if the environmental cleanliness level fluctuates several times higher than the class 10,000 level, the impact of air exposure during the installation of the LS-related system is controllable.
@@ VI.3 Structure outside the acrylic
The dust deposited on the outer surface of the acrylic sphere, PMTs, and SS truss during the construction process, can produce gamma rays due to its natural radioactivity. The gamma rays can penetrate through the water layer and the acrylic sphere and enter the LS, which contributes to the background. Based on the JUNO software framework Lin and others (2023), we simulated the natural radioactivity on the outer surface of the acrylic sphere, PMTs, and the SS truss Abusleme and others (2021). The probability for the deposited energy of being greater than 0.7 MeV and be positioned within the fiducial volume (R 17.2 m) of the LS for each natural radioactive nuclide is listed in Table 6. For 238U and 232Th, the full decay chain was simulated. Taking the acrylic outer surface as an example, 1 Bq of 238U will lead to 0.001 Hz event rate in the LS with deposited energy larger than 0.7 MeV at a position with radius smaller than 17.2 m, while the rate is one order of magnitude larger for 232Th, which has more energetic gamma rays in the chain (208Tl). At the same time, the radon produced by 238U in the dust will enter the shielding water, causing an increase of radon concentration.
| Acrylic outer surface | PMTsTruss | ||
| Radius [m] | 17.8 | 19.2-20.0 | |
| Probability | 238U | 1.110-3 | 3.410-6 |
| 232Th | 1.310-2 | 1.410-5 | |
As shown in Table 2, for flat plates made of PTFE and acrylic, the results for the upward-facing orientation are quite similar ( 1000 pg/d/m2). Although there are differences in the results for the downward-facing and side-facing orientations, their contributions are relatively minor ( 500 pg/d/m2) compared to the upward-facing results. A similar trend is observed for PFA bottles and flasks. This indicates that gravitational settling of dust during environmental exposure plays a dominant role. Therefore, even though the previous experiments did not simulate all detector material types, we can still refer to the results from PFA bottles and flasks when evaluating deposition rates. Since the aforementioned detector structures are uniformly spherical, we intend to use the measurement results from PFA flasks for background evaluation.
The installation situation and progress for acrylic sphere, PMTs and SS truss are summarized in Table 7, as well as the expected U/Th settling rate. Details on the construction can be found in Section IV.2. The exposure here only include the time after the final cleaning and before filling. The U/Th settling used in the calculation is the results of PFA flask from Table 3 during the exposure with the sphere facing upwards (No. 5 for acrylic and truss, No. 6 for PMTs layer). Based on the U/Th efficiency-to-singles-rate ratio within the FV for energies above 0.7 MeV from simulation (Table 6), the single-event rate originating from deposited U/Th on the detector surface inside the LS has been calculated and is presented in the penultimate row of the table. Taking into account the 9-kt UPW volume between the acrylic vessel and the truss, the rate of radon release from deposited U/Th of the detector surface to the water has also been calculated and is shown in the final row. Taking the acrylic column as an example, the total activity of 238U is 0.37 Bq. Assuming secular equilibrium, the activity of 222Rn is also 0.37 Bq. Dividing this by the volume of the inner water layer (9103 m3) gives a radon concentration of 0.04 mBq/m3. Overall, the U/Th contamination introduced during the installation of the detector due to environmental exposure contributes approximately 2 mHz to the single-event rate in the LS fiducial volume (R 17.2 m), accounting for about 0.03% of the total detector contribution (7 Hz Abusleme and others (2021)). The impact of radon released from the decay chain products of 238U deposited on the detector surface into the water is at a level of 2 mBq/m3 by assuming equilibrium rate as 238U, which is approximately 20% of the water radon target (10 mBq/m3). Additionally, some of the deposited dust can be further purified through the water circulation process.
| Acrylic | PMTs | SS truss | |
@@ VII Summary
The installation of the JUNO detector spanned three years. To ensure the detector’s cleanliness, the on-site installation environment was managed under clean-room conditions, with an average cleanliness level of class 74,000. After the installation of the acrylic sphere and before the filling with LS, the first step of the cleaning involved a water mist dust reduction inside the sphere, which improved the air cleanliness to a level of class 100 to 1,000. The second step involved using a 3D high-pressure nozzle to rinse the inner surface of the acrylic sphere, further enhancing the cleanliness of the surface in contact with the LS. It should be noted that both the purification and filling systems in contact with the LS and the inner surface of the acrylic vessel underwent rigorous cleaning and validation procedures to achieve the highest attainable level of surface cleanliness. However, the amount of residual impurities remaining on these surfaces and the extent to which they may contaminate the LS remain unknown and can only be evaluated through actual measurement data.
Surface deposition experiments for U/Th on various containers were conducted in the JUNO experimental hall. Based on experimental results, the total U/Th deposition on the acrylic vessel, PMTs, and SS truss during the detector installation process has been evaluated. Using detector simulations, the efficiency-to-singles-rate ratio for U/Th within the FV at energies above 0.7 MeV was found to be 0.1% for 238U and 1.3% for 232Th originating from the acrylic surfaces. In contrast, the corresponding efficiency for contributions from the PMTs and SS truss is approximately three orders of magnitude lower due to the presence of about 3 m of water shielding. Consequently, the single-event rate in the LS FV resulting from deposited surface dust accounts for only about 0.03% of the total rate (7 Hz) arising from all detector materials. The radon contamination level in the inner water region (approximately 9 kt) between the acrylic vessel and the SS truss, originating from U/Th deposits on the surfaces, is approximately 2 mBq/m3. This value meets only 20% of the maximum permitted limit of 10 mBq/m3.
@@ Acknowledgements
This work is supported by the National Key Research and Development Program of China (2024YFE0110503), the Youth Innovation Promotion Association of the Chinese Academy of Sciences, CAS Project for Young Scientists in Basic Research (YSBR-099).